# VersaTune: An Efficient Data Composition Framework for Training Multi-Capability LLMs

### v4

@@ 2.1. Transformer Models
2.2. LLMs Training
Large Language Models (LLMs) are Transformer-based models containing an extensive parameter count in the billions or greater, which are trained on a vast scale of datasets (Zhao et al., 2023b). The training process of LLMs generally involves the pretraining and fine-tuning stages. We have outlined several concepts about LLMs training.
2.2.1. Pretraining
Large Language Models (LLMs) establish basic knowledge abilities, including language understanding and text generation, during the pretraining stage (Brown et al., 2020). In this stage, LLMs engage in unsupervised training through the processing of extensive raw text corpora, thereby enhancing their capabilities in language modeling. For a given sequence , the typical task for LLMs involves the prediction of the subsequent token given the preceding tokens as contextual input. The goal is to maximize the likelihood function presented in Equation 1:
Beginning of Sequence (¡BOS¿) During the above process, the Beginning of Sequence (¡BOS¿) token plays an important role, which serves as a signal to the model that the input sequence is starting (Brown et al., 2020; Li et al., 2023a; Wu et al., 2025). It can be thought of as a special marker that indicates the start of a new sequence, allowing the model to reset its context and begin processing a new piece of text. In the context of pretraining, the ¡BOS¿ token is used to initialize the input to the model, and it can be concatenated with the actual text data to form the input sequence. This token helps the model to differentiate between the start of a new input and the continuation of an existing one. It is particularly crucial in tasks where the model needs to generate text or understand the beginning of a new sentence or document, which helps the model to learn the boundaries of text sequences and to better model the statistical properties of the language data it is trained on. The use of ¡BOS¿ tokens, along with other special tokens like ¡EOS¿ (End of Sequence), helps the model to learn the boundaries of text sequences and to better model the statistical properties of the language data it is trained on.
2.2.2. Supervised Fine-Tuning
@@
The Supervised Fine-Tuning (SFT) stage involves further training on the pretrained model to refine it task-solving capabilities and ensure greater alignment with human instructions across multiple domains (Zhao et al., 2023b; Nie et al., 2023). During the SFT phase, it is a typical practice to employ datasets specific to a particular domain for the fine-tuning of LLMs, which may lead to a decline in performance on non-target domains, a phenomenon commonly referred to as catastrophic forgetting. We conducted experiments on open-sourced models including LLaMA (Dubey et al., 2024; Touvron et al., 2023a, b) and Qwen (Bai et al., 2023; Yang et al., 2024) series to assess how the model’s proficiency in other domains changes when fine-tuned with data from a single domain, as illustrated in Table 1 and Figure 2. We have regulated the number of training instances per epoch to a fixed count of 10,000. More details about training and evaluation settings can be found in Section 5.1. Our findings indicate that when a model is trained exclusively with data from a single domain, its performance on tasks from other domains tends to degrade progressively over the course of training.
While recent research has delved into exploring fine-tuning methods for multi-task enhancement (Dong et al., 2023; Sanh et al., 2021), they are still in their early stages. However, as shown by proprietary models such as GPT-4 (Achiam et al., 2023) and Gemini (Team et al., 2023), which exhibit outstanding multi-task performance, improving a model’s versatile capabilities across various domains during the SFT phase is crucial. Therefore, our work systematically investigates methods to enhance multi-domain performance during the SFT stage to bridge this gap.
3. VersaTune
In this section, we introduce VersaTune, a novel data composition framework designed for multi-capability training. This framework aims to effectively compose data from multiple domains and optimize the data composition during training. Figure 3 presents the workflow of VersaTune, which generally contains two phases.
3.1. Phase 1: Knowledge Distribution Detection
3.1.1. Knowledge Consistency Training
Previous research on data mixing ratios during the SFT phase for LLMs has predominantly focused on enhancing capabilities within a specific domain, often utilizing only data from that domain or employing heuristic, experience-based data proportions. We argue that such data mixing strategies can significantly impair the LLM’s abilities in other domains. In the fine-tuning stage, maintaining a robust overall capability across various domains is crucial.
What data mixing strategy effectively boosts the versatile performance of LLMs across different domains during the SFT phase? We propose the following statement:
Statement 1 (Knowledge Consistency Training). An LLM fine-tuned with domain-specific data proportions that align with its pretrained output distributions will exhibit enhanced and balanced performance across these domains, compared to a model fine-tuned with a non-matching data distribution. Formally, the relationship can be represented as:
where denotes the set of all possible data points.
3.1.2. Domain Knowledge Distribution Detection
Drawing on prior research into knowledge identification methods (Gekhman et al., 2024; Zhao et al., 2023a) and training data inference strategies for LLMs (Ding et al., 2022; Hayase et al., 2024), we propose a structured approach to efficiently detect domain knowledge based on statistics. The process involves the generation of textual outputs from the base model waiting to be fine-tuning, followed by their classification into predefined domains referring to a proprietary model . The above process is repeated multiple times to ensure statistical robustness.
Assuming the data corpus encompasses distinct domains, as shown in Algorithm 1, we first prompt the base model poised for fine-tuning with the Beginning of Sequence (<BOS>) tokens to generate a set of data entries (Line 3). Subsequently, we employ a proprietary model to infer the probabilities that these entries belong to each domain (Line 5-7). We then calculate a weighted average of the probability distributions for all data across these domains, thereby deriving the domain knowledge distribution of the current base model (Line 9). To ensure statistical robustness, the process is iteratively conducted times, and we use the mean of these iterations as the estimated result for the domain knowledge distribution.
Input: Base model to be fine-tuned , Proprietary model , Hyperparameters: sample number , maximum iterations
Parameter: Data samples generated from the base model
Output: Domain distribution
Define : domain probability distribution of data sample
3.2. Phase 2: Fine-Tuning Multi-Ability LLMs Efficiently
Setting. Our goal is to construct a composite dataset covering specific domains, which can be denoted as with each tuple representing a specific domain and its corresponding proportion, such that training a model on dataset could achieve overall lower loss on a uniformly distributed composite target validation dataset or meet the flexible domain expansion while preserving the performance of other domains. The specialized capabilities of LLMs are ultimately measured using downstream tasks related to different domains (e.g., FinBen (Xie et al., 2024a) for financial performances).
3.2.1. Learnable Potential and Forgetting Degree of Knowledge
3.2.2. Effective Multi-Ability Fostering
Input: Base model to be fine-tuned , Domain reference loss , Hyperparameters: magnitude of adjustment , number of training steps
Parameter: Data proportion of the SFT dataset
Output: Fine-tuned multi-ability model
Define : learnable potential of the current domain
As detailed in Algorithm 2, in the pursuit of balanced domain expertise enhancement, we first blended the domain proportions in accordance with the base model’s intrinsic domain knowledge distribution detected by Algorithm 1 (Line 1). Then at each training step , we assigned a learnable potential score to each domain based on the methodology outlined in Equation 3. These scores were then utilized to fine-tune the representation of each domain within the composite SFT dataset, ensuring a balanced development of competencies across all domains throughout the fine-tuning process (Line 3-7). At the same time, the parameters of model are updated based on the gradients computed through backpropagation (Line 8). This adaptive approach is imperative to harmonize the progression of capabilities in different domains and to optimize the model’s performance on a multifaceted array of tasks.
3.2.3. Flexible Domain Expansion
@@
Input: Base model to be fine-tuned , Domains that require enhanced cultivation , Domain reference loss , Hyperparameters: number of training steps , magnitude of adjustment , extent of domain proportion adjustment , variation threshold
@@
When conducting fine-tuning on a pretrained model, there are instances where we aim to particularly enhance the model’s performance on specific domain tasks. Consequently, our algorithmic framework ought to possess the flexibility to accommodate domain expansion and generalize effectively. Building upon Statement 1, we present the following statement tailored for domain expansion:
Statement 2
where is the data distribution in the given domain , and is the adjustment factor.
where is the increased adjustment factor, and is the original ratio of other domain knowledge relative to .
As outlined in Algorithm 3, we initially establish the data distribution based on the knowledge detected from the original pretrained model (Line 1). At each training step , we calculate the learnable potential and forgetting degree scores for each domain (Line 4-5), and assign domain weights for the current training phase following the method from Algorithm 2 (Line 6). A trade-off is necessary between the remaining learning margin of the domain that requires focused cultivation and the model’s forgetting degree towards other non-target domains: if the improvement benefit of the specific domain exceeds the average forgetting degree of the other domains (ratio greater than ), we increase the data weight of the current specific domain by , and proportionally reduce the weights of the other non-target domains according to Equation 6 (Line 8-9); otherwise, we maintain the current domain distribution and only perform minor adjustments and renormalization as described in Algorithm 2 (Line 10-11). Subsequently, we update the parameters of the target model (Line 13).
4. Implementation
VersaTune offers a comprehensive solution for data composition aimed at enhancing performance across multiple domains. In this section, we will briefly discuss the implementation of key components and the setting of hyperparameters.
@@ 4. Implementation
Domain Knowledge Detection. During the knowledge distribution detection phase for our target models, we employed Qwen2.5-72B-Instruct111https://huggingface.co/Qwen/Qwen2.5-72B-Instruct as the proprietary model . For each target model slated for fine-tuning, we prompted the generation of data samples using the Beginning of Sequence () token, with the sample number set at . These samples were subsequently assessed by the proprietary model to ascertain their probabilistic affinity for several domains, including law, medicine, finance, science, code, and others. To ensure the reliability of our statistical outcomes, the entire process was iterated 5 times, with the maximum number of iterations set at . The average knowledge distribution was then computed across these iterations. Empirically, with a dataset of samples, the distribution of sequences generated by across domains demonstrated a high degree of consistency, with an overall variance not exceeding 1.874%. The final domain knowledge distribution for each open-source model is depicted in the stacked bar chart presented in Figure 4. The pre-existing domain knowledge distribution varies among different models. Therefore, it is essential to develop a data composition strategy that is tailored to the specific model being trained.
Hyper-Parameters Setting. During the multi-domain task fine-tuning of LLMs, we configured the number of training steps in Algorithm 2 and Algorithm 3 to 4 epochs. We experimented with various magnitude of adjustment, specifically , and observed consistent weight ordering across domains, which far outperformed our baselines (detailed in Section 5.2). Based on these experimental outcomes, we set the magnitude of adjustment to . Additionally, in the context of domain expansion, we set the increment for the target domain to per training step, considering the overall domain weight distribution across models. The variation threshold, denoted as , reflects the trade-off between enhancing specific domain skills and mitigating the loss of capabilities in non-target domains, where we assigned a weight of .
5. Experiments
@@
Lawyer-Instruct | 9241 | Reformatted from LawyerChat Dataset222https://huggingface.co/datasets/Alignment-Lab-AI/Lawyer-chat | 1 | |
@@
Code Alpaca | 20022 | Generate Based on Self-Instruct (Wang et al., 2022) | 1 | |
Alpaca | 49,087 | Generate Based on Self-Instruct (Wang et al., 2022) | 1 | Excluding Samples of Other Domains |
@@
Law | LegalBench (Guha et al., 2024) | 90,394 (164 sub-tasks) | English | 1 |
LawBench (Fei et al., 2023) | 10,000 (20 sub-tasks) | Chinese | 1 | |
Medicine | MedQA (Jin et al., 2020) | 1,273 | English | 1 |
MedMCQA (Pal et al., 2022) | 4,183 | English | 1 | |
Finance | FinEval (Zhang et al., 2023) | 4,661 (34 sub-tasks) | Chinese | 1 |
FinanceIQ (Zhang and Yang, 2023) | 7,173 (10 sub-tasks) | Chinese | 5 | |
Science | SciEval (Sun et al., 2024) | 15,901 | English | 1 |
MMLU-Sci (Hendrycks et al., 2020) | 2,999 (14 sub-tasks) | English | 0 | |
Code | HumanEval (Chen et al., 2021) | 164 | English | 0 |
MBPP (Austin et al., 2021) | 974 | English | 0 | |
Other (General) | AGIEval (Zhong et al., 2023) | 8,062 (20 sub-tasks) | English, Chinese | 0 |
HellaSwag (Zellers et al., 2019) | 10,003 | English | 0 |
In this section, we first describe the details of our experimental setup (Section 5.1), including training datasets, models used for experimentation and benchmarks. We then outline the baseline methods we use for comparison (Section 5.2) and our experimental results to prove the effectiveness of VersaTune (Section 5.3).
@@
Baichuan-2-7B | Uniform Distribution | 15.71 | 30.72 | 23.45 | 27.57 | 33.50 | 2.71 | 9.30 | 42.89 | 5.67 | 3.44 | 20.16 | 71.40 |
@@
Baichuan-2-13B | Uniform Distribution | 47.66 | 34.85 | 32.98 | 36.54 | 37.54 | 32.85 | 45.72 | 50.77 | 36.54 | 38.55 | 36.89 | 73.50 |
@@
VersaTune | 60.59↑ | 46.58↑ | 50.24↑ | 45.15↑ | 65.84↑ | 78.68↑ | 62.89↑ | 82.86↑ | 82.64↑ | 81.48↑ | 55.52↑ | 75.98↑ |
5.1. Experimental Setup
Training Datasets. We have collected datasets spanning 6 domains for supervised fine-tuning (SFT), including Lawyer-Instruct333https://huggingface.co/datasets/Alignment-Lab-AI/Lawyer-Instruct, the training portion of MedQA (Jin et al., 2020), Finance Alpaca444https://huggingface.co/datasets/gbharti/finance-alpaca, Sonnet3.5 Science Conversations555https://huggingface.co/datasets/jeffmeloy/sonnet3.5_science_conversations, Code Alpaca666https://github.com/sahil280114/codealpaca and Alpaca (Taori et al., 2023), denoted as , to represent SFT datasets with respect to law, medicine, finance, science, code as well as general capabilities. In order to prevent domain overlap, we curated the Alpaca dataset by excluding data pertaining to the other specific five domains, keeping only the general domain instances unrelated to them. Details of the training datasets can be found in Table 2.
Model Details. We use LLaMA (Dubey et al., 2024; Touvron et al., 2023a, b), Qwen (Bai et al., 2023; Yang et al., 2024) and Baichuan (Yang et al., 2023) series as our pretrained language models . For scenarios aimed at fostering multi-ability, we trained and assessed models including Baichuan-2-7B, LLaMA-3-8B, Baichuan-2-13B, Qwen-1.5-7B, Qwen-2-7B, Qwen-2.5-7B, and Qwen-2.5-14B. For domain expansion, the training and evaluations were performed using the Baichuan-2-7B and Baichuan-2-13B models. Regarding reference models, for the LLaMA series, we used the Sheared-LLaMA-1.3B (Xia et al., 2023) as a lightweight reference model; as for Qwen series, we utilized Qwen-2-1.5B and Qwen-2.5-1.5B as our reference models.
Training Details. All experiments were conducted based on full-parameter fine-tuning, during which we utilized a learning rate scheduler featuring linear warm-up and cosine decay, peaking at a learning rate of 2e-5, alongside a warmup ratio of 0.03, a weight decay of 0.0 and a batch size of 128 for 4 epochs. To maintain consistency, the total volume of training data across domains was controlled to 60,000 per epoch, with each domain’s samples being downsampled or upsampled according to the corresponding weights during the mixing process. We conducted all fine-tuning and evaluation experiments on NVIDIA RTX H800.
Benchmarks. We evaluate the performance of the models on downstream tasks across various domains, using two relevant benchmarks for each domain. Details of the datasets are provided in Table 3. Specifically, for the MedMCQA dataset, since the standard answers for the test set are not publicly available, we conducted our evaluations using the validation dataset. For the MMLU dataset, we selected 14 sub-tasks to construct the MMLU-Sci subset (Zhang et al., 2024) for testing, aiming to ensure a robust and thorough evaluation.
5.2. Baselines
We compare VersaTune with the following baselines. For the scenario of effective multi-ability fostering: (1) The simplest baseline is uniform distribution, where each domain has an equal weight proportion. (2) Inverse distribution assigns the proportionate weights to each domain in an inverse manner to the detected knowledge distribution. (3) VersaTune constant is implemented with fixed domain weights derived from the knowledge distribution obtained from probing the target model prior to fine-tuning, where we ablate the components of dynamic adaptation in Algorithm 2 for an in-depth analysis. Under the case of flexible domain expansion: (1) 100% specific domain strategy is a common practice to employ datasets consisting exclusively of data from a single domain during the fine-tuning stage. (2) Domain increase with uniform distribution of remainder elevates the proportion of a specific domain, while the remaining domains receive the balance of the distribution in an evenly distributed manner.
5.3. Results

### v5

@@ A.2 Analysis on Catastrophic Forgetting
During the SFT phase, it is a typical practice to employ datasets specific to a particular domain for the fine-tuning of LLMs, which may lead to a significant performance drop of knowledge in non-target domains, a phenomenon commonly referred to as Catastrophic Forgetting Kaushik et al. (2021); McCloskey and Cohen (1989). We conducted experiments on open-sourced models including LLaMA Dubey et al. (2024); Touvron et al. (2023a, b) and Qwen Bai et al. (2023); Yang et al. (2024) series to assess how the model’s proficiency in other domains changes when fine-tuned with data from a single domain, as depicted in Table 2 and Figure 6. We have regulated the number of training instances per epoch to a fixed count of 10,000. More details on training and evaluation settings can be found in Section 3.1 and Appendix C. Our findings indicate that when a model is trained exclusively with data from a single domain, its performance on tasks from other domains tends to degrade progressively over the course of training. This experimental outcome has provided significant motivation and direction for our work.
Appendix B Method Details
B.1 Algorithm for Flexible Domain Expansion
Here we provide the detailed algorithm of flexible domain expansion (Section 2.2.3). As outlined in Algorithm 3, we initially establish the data distribution based on the knowledge detected from the original pretrained model (Line 1). At each training step , we calculate the learnable potential and forgetting degree scores for each domain (Line 4-5), and assign domain weights for the current training phase following the method from Algorithm 2 (Line 6). A trade-off is necessary between the remaining learning margin of the domain that requires focused cultivation and the model’s forgetting degree towards other non-target domains. If the improvement benefit of the specific domain exceeds the average forgetting degree of the other domains (ratio greater than ), we increase the data weight of the current specific domain by , and proportionally reduce the weights of the other non-target domains according to Equation 5 (Line 8-9). Otherwise, we maintain the current domain distribution and only perform minor adjustments and renormalization as described in Algorithm 2 (Line 10-11). Subsequently, we update the parameters of the target model (Line 13).
@@ B.1 Algorithm for Flexible Domain Expansion
Input: Base model to be fine-tuned , Domains that require enhanced cultivation , Domain reference loss , Hyperparameters: number of training steps , magnitude of adjustment , extent of domain adjustment , variation threshold
@@
Appendix C Experiments Details
C.1 Knowledge Distribution Detection
@@
During the knowledge distribution detection phase for our target models, we have manually annotated 120 samples (20 samples for each domain) to fine-tune Qwen2.5-72B-Instruct555https://huggingface.co/Qwen/Qwen2.5-72B-Instruct, and employed the trained model as the proprietary model . For each target model slated for supervised fine-tuning, we prompted the generation of data samples using the Beginning of Sequence () token, with the sample number set at . These samples were subsequently assessed by the proprietary model to ascertain their probabilistic affinity for several domains, including law, medicine, finance, science, code, and others. To ensure the reliability of our statistical outcomes, the entire process was iterated 5 times, with the maximum number of iterations set at . The average knowledge distribution was then computed across these iterations. Empirically, with a dataset of samples, the distribution of sequences generated by across domains demonstrated a high degree of consistency, with an overall variance not exceeding 1.874%. The final domain knowledge distribution for each open-source model is depicted in the stacked bar chart presented in Figure 7. The pre-existing domain knowledge distribution varies among different models. Therefore, it is essential to develop a data composition strategy that is tailored to the specific model being trained.
C.2 Training Details
Models and Implementation. All experiments were conducted based on full-parameter fine-tuning, during which we utilized a learning rate scheduler featuring linear warm-up and cosine decay, peaking at a learning rate of 2e-5, alongside a warmup ratio of 0.03, a weight decay of 0.0 and a batch size of 128 for 4 epochs. For scenarios aimed at fostering multi-ability, we trained and assessed models including LLaMA-2-7B, LLaMA-3-8B, LLaMA-2-13B, Qwen-2-7B, Qwen-2.5-7B, Qwen-2.5-14B, and Qwen-2.5-32B. In the context of domain expansion, the training and evaluations were performed using the Qwen-2.5-7B and Qwen-2.5-14B models. The total number of samples per epoch was set to , with each domain’s samples being downsampled or upsampled according to the corresponding weights during the mixing process. Regarding reference models, for the LLaMA series, we used the Sheared-LLaMA-1.3B Xia et al. (2023) as a lightweight reference model; as for the Qwen series, we utilized Qwen-2-1.5B and Qwen-2.5-1.5B as our reference models.
Training Datasets. For training, we selected representative datasets for each domain, which exhibit significant differences in format, sentence length, and domain-specific content. These differences reflect the heterogeneity of training data across various domains during the fine-tuning stage. Further details about these datasets can be found in Table 3. Specifically, for the Alpaca dataset, which we utilize for representing the general domain, we have excluded data related to law, medicine, finance, science, and code domains to ensure the precision and authenticity of the actual domain weight.
@@ C.2 Training Details
Lawyer-Instruct | 9241 | Reformatted from LawyerChat Dataset666https://huggingface.co/datasets/Alignment-Lab-AI/Lawyer-chat | 1 | |
@@
Code Alpaca | 20022 | Generate Based on Self-Instruct Wang et al. (2022) | 1 | |
Alpaca | 49,087 | Generate Based on Self-Instruct Wang et al. (2022) | 1 | Excluding Samples of Other Domains |
C.2.1 Hyper-Parameters Setting
During the multi-domain task fine-tuning of LLMs, we configured the number of training steps in Algorithm 2 and Algorithm 3 to 4 epochs. We experimented with various magnitude of adjustment, specifically , and observed consistent weight ordering across domains, which far outperformed our baselines (detailed in Section 3.2). Based on these experimental outcomes, we set the magnitude of adjustment to . Additionally, in the context of domain expansion, we set the increment for the target domain to per training step, considering the overall domain weight distribution across models. The variation threshold, denoted as , reflects the trade-off between enhancing specific domain skills and mitigating the loss of capabilities in non-target domains, where we assigned a weight of .
C.3 Evaluation Details
We evaluate the performance of the models on downstream tasks across various domains, using two relevant benchmarks for each domain. Details of the datasets are provided in Table 4. Specifically, for the MedMCQA dataset, since the standard answers for the test set are not publicly available, we conducted our evaluations using the validation dataset. For the MMLU dataset, we selected 14 sub-tasks to construct the MMLU-Sci subset Zhang et al. (2024) for testing, aiming to ensure a robust and thorough evaluation.
@@ C.3 Evaluation Details
Law | LegalBench Guha et al. (2024) | 90,394 (164 sub-tasks) | English | 1 |
LawBench Fei et al. (2023) | 10,000 (20 sub-tasks) | Chinese | 1 | |
Medicine | MedQA Jin et al. (2020) | 1,273 | English | 1 |
MedMCQA Pal et al. (2022) | 4,183 | English | 1 | |
Finance | FinEval Zhang et al. (2023) | 4,661 (34 sub-tasks) | Chinese | 1 |
FinanceIQ Zhang and Yang (2023) | 7,173 (10 sub-tasks) | Chinese | 5 | |
Science | SciEval Sun et al. (2024) | 15,901 | English | 1 |
MMLU-Sci Hendrycks et al. (2020) | 2,999 (14 sub-tasks) | English | 0 | |
Code | HumanEval Chen et al. (2021) | 164 | English | 0 |
MBPP Austin et al. (2021) | 974 | English | 0 | |
Other (General) | AGIEval Zhong et al. (2023) | 8,062 (20 sub-tasks) | English, Chinese | 0 |
HellaSwag Zellers et al. (2019) | 10,003 | English | 0 |
Appendix D More Experiment Results
D.1 Multi-Ability Fostering
We present the results of the ablation study along with the raw scores from domain-specific benchmarks. As shown in Table 5 and Figure 8, VersaTune constant is implemented with fixed domain weights derived from the knowledge distribution obtained by probing the target model prior to fine-tuning, where we ablate the components of dynamic adaptation in Algorithm 2 for an in-depth analysis. Additionally, to demonstrate the robustness of the dynamic adjustment more clearly, we compare the domain-level averaged performance of the VersaTune constant and VersaTune strategies, as depicted in Table 6.
@@ D.1 Multi-Ability Fostering
LLaMA-2-7B | Uniform Distribution | 15.71 | 30.72 | 23.45 | 27.57 | 33.50 | 2.71 | 9.30 | 42.89 | 5.67 | 3.44 | 20.16 | 71.40 |
@@
LLaMA-2-13B | Uniform Distribution | 47.66 | 34.85 | 32.98 | 36.54 | 37.54 | 32.85 | 45.72 | 50.77 | 36.54 | 38.55 | 36.89 | 73.50 |
@@
VersaTune | 60.59↑ | 46.58↑ | 50.24↑ | 45.15↑ | 65.84↑ | 78.68↑ | 62.89↑ | 82.86↑ | 82.64↑ | 81.48↑ | 55.52↑ | 75.98↑ | |
Qwen-2.5-32B | Uniform Distribution | 68.86 | 45.28 | 72.34 | 68.18 | 68.03 | 75.14 | 58.30 | 80.17 | 78.59 | 71.04 | 75.26 | 84.40 |
Inverse Distribution | 62.93↓ | 42.05↓ | 68.80↓ | 66.09↓ | 66.80↓ | 73.93↓ | 52.94↓ | 79.31↓ | 74.44↓ | 70.71↓ | 75.00↓ | 83.80↓ | |
VersaTune Constant | 71.98↑ | 54.93↑ | 75.42↑ | 71.05↑ | 69.38↑ | 77.07↑ | 65.87↑ | 82.11↑ | 82.38↑ | 77.16↑ | 74.97↓ | 84.82↑ | |
VersaTune | 75.67↑ | 56.76↑ | 78.72↑ | 72.36↑ | 70.50↑ | 78.80↑ | 70.77↑ | 85.23↑ | 86.60↑ | 79.89↑ | 75.81↑ | 84.75↑ |
Model | Method | Law | Medical | Finance | Science | Code | General | Avg. |
LLaMA-2-7B | VersaTune Constant | 28.51↓ | 33.45↓ | 21.09↓ | 47.89↓ | 13.47↓ | 47.10↓ | 31.92↓ |
VersaTune | 29.75 | 37.90 | 32.66 | 53.41 | 15.65 | 48.22 | 36.27 | |
Qwen-2-7B | VersaTune Constant | 39.29↓ | 40.01↓ | 50.73↓ | 56.61↓ | 51.29↓ | 60.57↓ | 49.75↓ |
VersaTune | 43.05 | 43.36 | 64.67 | 61.00 | 52.90 | 60.85 | 54.31 | |
Qwen-2.5-7B | VersaTune Constant | 41.99↓ | 39.96↓ | 55.71↓ | 64.00↓ | 55.63↓ | 60.82↓ | 53.02↓ |
VersaTune | 44.20 | 43.19 | 65.79 | 71.15 | 60.97 | 61.82 | 57.85 | |
LLaMA-3-8B | VersaTune Constant | 42.80↓ | 33.74↓ | 31.60↓ | 55.41↓ | 31.55↓ | 49.48↓ | 40.76↓ |
VersaTune | 43.77 | 43.97 | 41.19 | 58.64 | 32.28 | 51.20 | 45.18 | |
LLaMA-2-13B | VersaTune Constant | 46.26↓ | 41.72↓ | 38.90↓ | 56.12↓ | 42.75↓ | 56.09↓ | 46.97↓ |
VersaTune | 48.01 | 46.73 | 47.30 | 63.36 | 46.15 | 57.14 | 51.45 | |
Qwen-2.5-14B | VersaTune Constant | 51.29↓ | 44.88↓ | 34.14↓ | 63.39↓ | 78.53↓ | 65.75↓ | 56.33↓ |
VersaTune | 53.59 | 47.70 | 72.26 | 72.88 | 82.06 | 65.75 | 65.71 | |
Qwen-2.5-32B | VersaTune Constant | 63.46↓ | 73.24↓ | 73.23↓ | 73.99↓ | 79.77↓ | 79.90↓ | 73.93↓ |
VersaTune | 66.22 | 75.54 | 74.65 | 78.00 | 83.25 | 80.28 | 76.32 |
Comparison with Frontier Models. Furthermore, to demonstrate the efficacy of VersaTune across diverse domain tasks, we conducted a comparative analysis between Qwen-2.5-32B + VersaTune and frontier models across various domain-specific tasks, with results visualized in Figure 9. Such experimental results indicate that Qwen-2.5-32B equipped with VersaTune enhances multi-domain performance to the state-of-the-art levels, which even outperforms frontier models like GPT-4o, Claude3.5-Sonnet and DeepSeek-V3 by 0.86%, 4.76% and 4.60% on the overall domain capabilities, respectively. This comparison underscores the superior performance of our VersaTune in advancing multi-domain capabilities.
D.2 Flexible Domain Expansion